OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-3, 5) and (-2, -6)

OpenStudy (anonymous):

First, you have to get the slope of the line, using these two points: \[\frac{ y _{2} - y _{1} }{ x _{2} - x _{1} } = m\] where "m" is the slope and (x1, y1) is (-3, 5) and (x2, y2) is (-2, -6).

OpenStudy (anonymous):

So whats the line?

OpenStudy (anonymous):

Once you calculate the slope, you can get the equation (in point-slope form) of the line as:\[y- y _{1} = m(x - x _{1})\] where you can use either of the points for (x1, y1). Rearranging the equation into the form ax + by = c will give you the standard form of the line. What did you calculate for the slope? Show all work, substitutions and simplifications.